A sample of 160 people has a mean age of 27 with a population standard deviation of 5. Test the hypothesis that the population mean is 26.7 at =0.05

1
Expert's answer
2022-06-09T14:47:35-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=26.7"

"H_1:\mu\not=26.7"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

The rejection region for this two-tailed test is "R = \{z:|z|> 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{27-26.7}{5/\sqrt{160}}=0.759"


6. Since it is observed that "|z|=0.759<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed is "p=2P(Z>0.759)=0.447853," and since "p=0.447853>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 26.7, at the "\alpha = 0.05" significance level.